Hausmitteilung

An alle Mitarbeiterinnen und Mitarbeiter Abteilung Kommunikation und Datenverarbeitung

Leiter

Dr. habil. R. Fülle

Telefon (03 51) 2 60-32 64 Telefax (03 51) 2 60-20 48 fuelle@fz-rossendorf.de

16. Januar 2001

Betreff ...

Sehr geehrte ...

Text

signname